# Target Detection in Sea Clutter with Application to Spaceborne SAR Imaging

### v3

@@ Abstract
In this paper, the challenging task of target detection in sea clutter is addressed. We analyze the statistical properties of the signals which have been received from the scene and based on that, we model the amplitude of the signals that have been reflected from the sea clutter according to several well-known probability distributions. Next, by exploiting the Kullback-Leibler (KL) divergence metric as a goodness-of-fit test, we will demonstrate that among the proposed probability distributions, the Weibull distribution can model the statistical properties of the sea clutter with higher accuracy. Subsequently, we utilize the aforementioned information to design an adaptive threshold based on the Constant False Alarm Rate (CFAR) algorithm to detect the energy of the targets which have been buried in the sea clutter.
Thorough analysis of the experimental data gathered from the Canadian RADARSAT-1 satellite demonstrates the overall effectiveness of the proposed method.
Index Terms:
Target detection, sea clutter, Kullback-Leibler divergence metric, Weibull distribution, CFAR.
@@ I Introduction
As an all-weather, day-night, and active device, Synthetic Aperture Radar (SAR) is capable of providing crucial information from the surface of the Earth [1, 3, 4, 5, 2, 6]. Following the image reconstruction, post processing procedure is performed to obtain more critical information from the scene. Target detection is one of the areas that requires further analysis of the reconstructed images. In this paper, we specifically focus on the targets buried in sea clutter. The goal is to establish a mechanism to extract the energy of the targets from sea clutter. Several techniques have been reported in the literature, such as the wavelet-based analysis [7] and fixed threshold-based method [8]. The problem with using a fixed threshold to detect targets from clutter is that some of the key features of the desired targets can be ignored while the undesired signals are retained. Moreover, methods such as wavelet-based analysis are not able to capture the statistical properties of the signals. This, in turn, hinders their ability to obtain all the critical information about the desired targets.
In this paper, we utilize an approach based on the Constant False Alarm Rate (CFAR) method [9, 10] which creates an adaptive threshold to extract the energy of the desired targets while suppressing the energy of the undesired signals. However, in order to be able to implement the CFAR-based adaptive threshold, we need to obtain the information about the statistical properties of the undesired signals which in this case is the sea clutter. To tackle this issue, we study several well-known probability distributions, namely, Weibull, Log-normal, Inverse Gaussian, Gamma, and Rayleigh and demonstrate that the Weibull distribution is a more appropriate distribution to model the statistical properties of the sea clutter in this paper.
@@
In [11, 12, 13], the authors have studied the statistical properties of the reflected SAR signals. In this paper, however, we specifically focus on sea clutter and also we present more statistical models which distinguishes our work from [11, 12]. We further describe the SAR imaging procedure in detail and present the experimental results to verify the effectiveness and accuracy of the proposed models which makes the material presented in this paper distinct from [11, 12, 13]. In contrast to [13], we will not consider the K distribution. The reason for this is that, the K distribution is formed by compounding two separate probability distributions, one representing the radar cross-section, and the other representing the speckle noise. However, through experimental results, we demonstrate that the Weibull distribution can model the sea clutter highly accurately. As a result, the K distribution is no longer considered as a possible candidate to model the statistical properties of the sea clutter in this paper.
@@
We present several statistical models to describe the sea clutter and at the end demonstrate that the Weibull distribution is the most appropriate distribution to model the statistical properties of the sea clutter. In order to determine which probability distribution can model the background sea clutter with higher accuracy, we exploit the Kullback Leibler (KL) divergence metric as a goodness-of-fit test. We then utilize the Weibull distribution and establish an adaptive threshold to extract the energy of the desired targets from sea clutter. The verification of the proposed approach is based on the real data gathered from the Canadian RADARSAT-1 SAR satellite [14]. The data has been gathered from English Bay in Vancouver Canada in which several ships have been located inside water which creates a perfect scenario to evaluate the results.
@@ II Model Description and Image Formation
Fig. 1 illustrates the system model in which the satellite is moving along its track and is collecting data from the surface of the Earth. The data collection is based on the strip-map mode.
[…]
Figure 1: The geometry of the model which shows a satellite-based SAR system in strip-map mode with non-zero squint angle.
The signal transmitted by the radar is a chirp signal which after being reflected from a point target and upon down-conversion is described as {dmath} s(t, η) = σ  w_r(t - 2R(η)c)w_a(η- η_c)×e^-j 4 πfcR(η)c + j πβt2 + j 4 πβt R(η)c, where $f_{c}$ is the carrier frequency and the parameter $\beta$ is given as $b/T$, in which $b$ and $T$ stand for the bandwidth and the chirp time, respectively. In addition, $w_{r}$ is a rectangular window with length $T$ and $t$ is referred to as fast-time parameter. In addition, $\sigma$ is the complex radar cross section for the point reflector, $\eta$ is referred to as slow-time parameter and $R(\eta)$ is the instantaneous radial distance between the radar and the target which is given as $R(\eta)=\sqrt{R^{2}_{0}+v^{2}(\eta-\eta_{c})^{2}}$. Moreover, $w_{a}$ is a rectangular window with its length equal to the synthetic aperture length divided by $v$, where $v$ is the speed of the satellite. The ultimate goal in SAR imaging is to estimate the complex radar section $\sigma$.
The Range-Doppler algorithm is the first algorithm that was developed for spaceborne SAR image reconstruction [15, 16, 17]. The first step in the Range-Doppler algorithm is to perform range compression which results in [3]
@@
where $p_{r}(x)=\frac{sin(\pi x)}{\pi x}$.
Next, we should compensate for the Range Cell Migration (RCM) phenomenon which is the result of coupling between the range and azimuth directions that is created because of the non-zero squint angle for the antenna. Consequently, after RCM compensation, we can describe the signal in the Doppler domain as [3] {dmath} S_rcmc(t, f_η) = σ  p_r(t - 2R0c)W_a(f_η - f_η_c) ×e^-j 4 πfcR0c e^j πf2ηKa, in which $K_{a}=\frac{2v^{2}}{\lambda R_{0}}$. The final step for image formation in the Range-Doppler algorithm, is to compress the data in the azimuth direction which upon performing this stage the compressed signal in the range and azimuth directions is given as [3] {dmath} s_rcac(t, η) = σ  p_r(t - 2R0c)p_a(η) e^-j 4 πfcR0c ×e^j 2 πfηc η, where $f_{\eta_{c}}$ is the Doppler centroid frequency [3, 19, 20, 18, 21].
@@ III Statistical Modeling
In this section, we attempt to model the amplitude of the reflected signal from sea clutter. The main candidate is the Weibull distribution which its probability density function is described as [22]
@@
Another possible option for consideration to model the sea clutter is the Log-normal distribution which its probability density function is given as [22]
@@
The next distribution that we consider in this paper, is the Inverse Gaussian distribution with the following probability density function [22]
@@
The Gamma distribution with 2 degrees of freedom is the next best choice to model the sea clutter. The probability density function for the Gamma distribution is expressed as [22]
@@
The last distribution that we study in this paper, is the Rayleigh distribution. The probability density function for the Rayleigh distribution is described as [22]
$p_{R}(x|\sigma)=\frac{x}{\sigma^{2}}e^{\displaystyle-\frac{x^{2}}{2\sigma^{2}}},\;\;\;x\geq 0,\;\;\sigma>0.$ (6)
Based on the given probability of false alarm ($p_{\rm fa}$), we can calculate the adaptive threshold $T_{a}$ as [9, 10]
[…]
where $X$ describes the statistics of the cell under test and $T_{a}$ is the adaptive threshold. Furthermore, $H_{0}$ represents the null hypothesis.
By using the experimental data that we present in this paper, we will demonstrate that among the proposed distributions, the Weibull distribution is the best distribution to model the sea clutter. Therefore, the calculation of the adaptive threshold is only performed for the Weibull distribution. Consequently, based on (7), the adaptive threshold $T_{\rm aW}$ for the Weibull distribution, which has been given in (2), is calculated as
@@
in which $x_{i}$ is the $i^{\rm th}$ realization of the random variable $x$.
In this paper, we implement the 2D CFAR [9, 10] algorithm. The implementation of the 2D CFAR method is based on the structure which has been shown in Fig. 2.
Refer to captionTraining CellsGaurd CellsCell Under Test
Figure 2: The structure of the 2D CFAR algorithm.
Consequently, the decision making process for the cell under test is expressed as [25, 12]
@@
where $X_{\rm CUT}$ is the amplitude of the cell under test. Furthermore, $\mu_{c}$ and $\sigma_{c}$ are the sample mean and standard deviation computed from the clutter data of the local background. Moreover, $Q=\frac{T_{\rm aW}}{\hat{\mu}}$ is the detector design parameter which defines the $p_{\rm fa}$ and is set empirically in which $T_{\rm aW}$ is the adaptive threshold which has been described in (8) and $\hat{\mu}$ denotes the mean value estimated from the underlying clutter model. In addition, the $H_{0}$ and $H_{1}$ are the null and alternative hypotheses, respectively [24, 10].
The null hypothesis, $H_{0}$, represents the case in which the cell under test has been occupied by clutter only and the alternative hypothesis, $H_{1}$, on the other hand, is for the case in which the target is present and the cell under test contains the energy of the target.
In the next section, we present the experimental results.
@@
In this section, we present the result of SAR image reconstruction based on experimental data gathered from the Canadian RADARSAT-1 satellite. The data is from Vancouver Canada. The specifications of the RADARSAT-1 satellite have been given in Table.I.
@@
FM rate of radar pulse ($\rm MHz/\mu s$) $\rm\beta$ $0.72135$
Chirp duration ($\rm\mu s$) $\rm T$ $41.75$
@@
We select the data related to English Bay. The reason is that, there are several ships in this scene that play the role of isolated strong reflectors which will assist to observe the effect of the RCM clearly and will also be utilized to perform Doppler centroid frequency estimation. Furthermore, the ships in the sea water are the main subject of the paper. In other words, we attempt to detect the ships in sea clutter which is the main goal of the paper.
In order to reconstruct the image, we apply the Range-Doppler algorithm to the raw data. Fig. 3-(a) shows the range compressed data based on (1). The range compressed energy of the ships can be seen as several skewed vertical lines. The skew demonstrates the effect of the RCM. In order to perform the RCM compensation we need to estimate the unambiguous Doppler centroid frequency.
@@
When the Doppler centroid frequency is larger than the PRF of the RADAR, there will be an ambiguity in the Doppler centroid frequency estimation. One way to estimate the unambiguous value for the Doppler centroid frequency is by analyzing the trajectory of the strong isolated targets. Fig. 3 shows the range compressed image of several ships in sea water. The skew in the trajectory of these targets is due to the non-zero Doppler centroid frequency. The energy of each of these targets is spread over several different range cells. The slope can easily be calculated and is equal to $0.034$ range samples per azimuth samples. To estimate the Doppler centroid frequency, the first step is to multiply the slope by $\rm\frac{c}{2Fr}$ to convert the slope from range samples to range distance and then multiply it by $\rm PRF$ to convert it from azimuth samples to azimuth time. Hence, we have $\frac{dR(\eta)}{d\eta}=198.23\;\rm m/s$. As a result, from $f_{\rm dc}=-\frac{2}{\lambda}\frac{dR(\eta)}{d\eta}$ [3, 19, 20, 18, 21], the Doppler centroid frequency can be calculated as $f_{\rm dc}=-7009\;\rm Hz$.
@@
Next, we estimate the fractional part of the Doppler centroid frequency which is $f^{\prime}_{\rm dc}$ [3, 19, 20, 18, 21]. The fractional part is essential in focusing the energy of the targets in the azimuth direction. Fig. 4 illustrates the power spectrum of the data in azimuth direction versus slow-time frequency component $f_{\eta}$. In order to reduce the effect of the noise, we have performed an averaging over the power spectrum of $2048$ range cells. In order to be able to calculate the frequency component at which the signal reaches its maximum value, we have performed curve fitting. From Fig. 4, we can estimate the fractional part of the Doppler centroid frequency as $f^{\prime}_{\rm dc}=531\;\rm Hz$. Finally, at the last stage, we perform the azimuth localization based on (1). Consequently, the reconstructed image based on the Range-Doppler algorithm is obtained which has been presented in Fig. 5.
@@
However, the image presented in Fig. 5 suffers from speckle noise [3]. In order to remove the effect of the speckle noise, we introduce a 2D $m\times n$ filter and slide it over the reconstructed image while solving the following optimization problem,
@@
where $a_{\rm ij}$ is the value for the $(\rm ij)^{\rm th}$ pixel and $a$ is the value selected by the optimization problem for the $(\lfloor\frac{n-1}{2}\rfloor+1\;\lfloor\frac{m-1}{2}\rfloor+1)^{\rm th}$ pixel. In fact, by solving the optimization problem in (11) we are filtering the image by the median filter. In other words, we replace the value of each pixel with the median of the neighboring pixels. This is proved to be a powerful method to decrease the effect of the speckle noise which, as a result, the fine structures of the image can be revealed. It should be noted that, the common practice to alleviate the effect of the speckle noise in SAR images is multi-look processing [3]. However, multi-look processing sacrifices the azimuth resolution in order to remove the speckle noise. In contrast, the median filtering approach, which we have proposed, can efficiently remove the speckle noise while it leaves the azimuth resolution intact. Furthermore, the median filtering does not smear the edges in the image which is considered to be a significant property.
@@
Fig. 6 shows the result of speckle noise reduction for the reconstructed image depicted in Fig. 5. To remove the effect of the speckle noise, we have applied the filter, given in (11), with $m=n=6$.
In order to analyze the effect of the sea clutter on target detection, we focus on the specific part of the reconstructed image shown in Fig. 5 which contains several ships in sea water. The related part of the data-set has been highlighted as the area inside the green rectangle in Fig. 5. The area inside the red rectangle contains the energy from sea water and is utilized to analyze the statistical properties of the sea clutter. Fig. 7 shows the histogram for part of the image related to the area which has been located inside the red rectangle. The estimated probability density functions for the Weibull, Log-normal, Inverse Gaussian, Gamma, and Rayleigh distributions are based on (2), (3), (4), (5), and (6), respectively.
@@
Figure 7: The histogram of the experimental data inside the red box which is related to the sea clutter as well as the estimated probability density functions based on (2), (3), (4), (5), and (6). The $I$ and $Q$ represent the imaginary and quadratic part of the estimated value for the complex radar cross section.
The estimated parameters for the given probability density functions have been presented in Table.II.
@@
From Table.III, it is evident that compared to the other probability distributions, the Weibull distribution, with the lowest KL distance, can model the statistical properties of the sea clutter with higher accuracy. Therefore, we only consider the Weibull distribution for the CFAR-based analysis of target detection in sea clutter.
Fig. 8 illustrates the result of applying the 2D CFAR algorithm to the part of the data-set inside the green rectangle in Fig. 5. The result presented in Fig. 8-(b) is remarkable in a sense that, by modeling the sea clutter based on the Weibull distribution and by exploiting the 2D CFAR technique we have been able to detect the ships in sea clutter with high accuracy.
Refer to captionAlong-Track [Samples]Slant-Range [Samples](a)
Refer to captionAlong-Track [Samples]Slant-Range [Samples](b)
@@
To achieve the result which has been displayed in Fig. 8-(b), we have set the number of guard cells for the upper and lower wings to 60 and for the left and right wings to 90. We have also selected the number of training cells for the upper, lower, left, and right wings to be equal to 5. The probability of false alarm has been selected as $p_{\rm fa}=10^{-6}$. The Weibull parameters are estimated as $\alpha=1.9521$ and $\beta=0.4835$. Consequently, based on (8), the adaptive threshold is calculated as $T_{\rm aW}=1.856$. Moreover, the sample mean and standard deviation for the background clutter, which have been described in (10), are estimated as $\mu_{c}=0.4286$ and $\sigma_{c}=0.2294$.
We consider another case in which one of the ships, as has been highlighted as the area inside the yellow rectangle in Fig. 5, is buried in the sea clutter. Moreover, it should be noted that, the image presented in Fig. 5 is contaminated with the speckle noise.
In Fig. 9, we have presented the reconstructed image from a single ship which has been highlighted as the area inside the yellow box in Fig. 5.
Refer to captionAlong-Track [Samples]Slant-Range [Samples](a)
Refer to captionAlong-Track [Samples]Slant-Range [Samples](b)
@@
The part of the image, shown in Fig. 9-(b), which has been highlighted as the area inside the red rectangle is utilized to estimate the histogram of the sea clutter and the result has been illustrated in Fig. 10.
@@
Figure 10: The histogram of the part of the reconstructed image shown in Fig. 9-(b) inside the red rectangle as a representation of the statistical properties of the sea clutter. The estimated probability density function is based on the Weibull Distribution.
As it is clear from Fig. 10, the Weibull distribution can model the histogram with high accuracy.
Refer to captionAlong-Track [Samples]Slant-Range [Samples]
Figure 11: The result of applying the 2D CFAR method to the image of the ship in sea clutter which has been illustrated in Fig. 9-(b).
Next, we apply the 2D CFAR technique to the reconstructed image presented in Fig. 9-(b). We have presented this remarkable result in Fig. 11. To achieve the result presented in Fig. 11, we have set the number of guard cells for the upper and lower wings to 35 and for the left and right wings to 60. We have also selected the number of training cells for the upper, lower, left, and right wings to be equal to 8. The probability of false alarm has been set to $p_{\rm fa}=10^{-6}$. The Weibull parameters are estimated as $\alpha=1.9912$ and $\beta=0.2841$. Consequently, based on (8), the adaptive threshold is calculated as $T_{\rm aW}=1.0621$. Furthermore, the sample mean and standard deviation for the background clutter, which have been utilized in (10), are estimated as $\mu_{c}=0.2518$ and $\sigma_{c}=0.1321$.

### v4

@@ Abstract
In this paper, the challenging task of target detection in sea clutter is addressed. We analyze the statistical properties of the signals which have been received from the scene and based on that, we model the amplitude of the signals that have been reflected from the background sea clutter according to several well-known probability distribution functions. Next, by exploiting the Kullback-Leibler (KL) divergence metric as a goodness-of-fit test, we will demonstrate that among the proposed probability distributions, the Weibull distribution can model the statistical properties of the background sea clutter with higher accuracy. Subsequently, we utilize the aforementioned information to design an adaptive threshold based on the Constant False Alarm Rate (CFAR) algorithm to detect the energy of the targets which have been buried in the sea clutter. Thorough analysis of the experimental data gathered from the Canadian RADARSAT-1 satellite demonstrates the overall effectiveness of the proposed method.
This paper has been published in the 2025 International Conference on Radar, Antenna, Microwave, Electronics, and Telecommunications (ICRAMET). ©IEEE
@@ I Introduction
As an all-weather, day-night, and active device, Synthetic Aperture Radar (SAR) is capable of providing crucial information from the surface of the Earth [1]. Following the image reconstruction, post processing procedure is performed to obtain more critical information from the scene. Target detection is one of the areas that requires further analysis of the reconstructed images. In this paper, we specifically focus on the targets buried in sea clutter. The goal is to establish a mechanism to extract the energy of the targets from sea clutter. Several techniques have been reported in the literature, such as the wavelet-based analysis [2] and fixed threshold-based method [3]. The problem with using a fixed threshold to detect targets from clutter is that some of the key features of the desired targets can be ignored while the undesired signals are retained. Moreover, methods such as wavelet-based analysis are not able to capture the statistical properties of the signals. This, in turn, hinders their ability to obtain all the critical information about the desired targets.
@@
In this paper, we utilize an approach based on the Constant False Alarm Rate (CFAR) method [4] which creates an adaptive threshold to extract the energy of the desired targets while suppressing the energy of the undesired signals. However, in order to be able to implement the CFAR-based adaptive threshold, we need to obtain the information about the statistical properties of the undesired signals which in this case is the sea clutter. To tackle this issue, we study several well-known probability density functions, namely, Weibull, Log-normal, Inverse Gaussian, Gamma, and Rayleigh and demonstrate that the Weibull distribution is a more appropriate distribution to model the statistical properties of the sea clutter. In [5, 6, 7, 8], the authors have studied the statistical properties of the reflected SAR signals. In this paper, however, we specifically focus on sea clutter and also we present more statistical models which distinguishes our work from [5, 6, 7]. We further describe the SAR imaging procedure in detail and present the experimental results to verify the effectiveness and accuracy of the proposed models which makes the material presented in this paper distinct from [5, 6, 7, 8]. In contrast to [8], we will not consider the K distribution. The reason for this is that, the K distribution is formed by compounding two separate probability distributions, one representing the radar cross-section, and the other representing the speckle noise. However, through experimental results, we demonstrate that the Weibull distribution can model the sea clutter highly accurately. As a result, the K distribution is no longer considered as a possible candidate to model the statistical properties of the sea clutter in this paper. In fact, after removing the effect of the speckle noise the K distribution is no longer capable of modeling the sea clutter accurately.
@@
Several statistical models are presented to describe the background sea clutter and at the end we demonstrate that the Weibull distribution is the most appropriate distribution to model the statistical properties of the background sea clutter. In order to determine which probability distribution can model the background sea clutter with higher accuracy, we exploit the Kullback Leibler (KL) divergence metric as a goodness-of-fit test. We then utilize the Weibull distribution and establish an adaptive threshold to extract the energy of the desired targets from the background sea clutter. The verification of the proposed approach is based on the real data gathered from the Canadian RADARSAT-1 SAR satellite [9]. The data has been gathered from English Bay in Vancouver Canada in which several ships have been located inside water which creates a perfect scenario to evaluate the results.
@@ II Model Description And Image Formation
The signal transmitted by the radar is a chirp signal which after being reflected from a point target and upon down-conversion is described as
$s(t,\eta)=\sigma\;w_{r}(t-\frac{2R(\eta)}{c})w_{a}(\eta-\eta_{c})\times e^{\displaystyle-j4\pi f_{c}\frac{R(\eta)}{c}+j\pi\beta t^{2}+j4\pi\beta t\frac{R(\eta)}{c}},$ (1)
where $f_{c}$ is the carrier frequency and the parameter $\beta$ is given as $b/T$, in which $b$ and $T$ stand for the bandwidth and the chirp time, respectively. In addition, $w_{r}$ is a rectangular window with length $T$ and $t$ is referred to as fast-time parameter. In addition, $\sigma$ is the complex radar cross section for the point reflector, $\eta$ is referred to as slow-time parameter and $R(\eta)$ is the instantaneous radial distance between the radar and the target which is given as $R(\eta)=\sqrt{R^{2}_{0}+v^{2}(\eta-\eta_{c})^{2}}$. Moreover, $w_{a}$ is a rectangular window with its length equal to the synthetic aperture length divided by $v$, where $v$ is the speed of the satellite. The Range-Doppler algorithm is the first algorithm that was developed for spaceborne SAR image reconstruction [10]. The first step in the Range-Doppler algorithm is to perform range compression which results in [1]
@@
where $p_{r}(x)=\frac{sin(\pi x)}{\pi x}$. Next, we should compensate for the Range Cell Migration (RCM) phenomenon which is the result of coupling between the range and azimuth directions that is created because of the non-zero squint angle for the antenna. Consequently, following the RCM compensation, the signal in the Doppler domain is described as [1]
$S_{\rm rcmc}(t,f_{\eta})=\sigma\;p_{r}(t-\frac{2R_{0}}{c})W_{a}(f_{\eta}-f_{\eta_{c}})\times e^{\displaystyle-j\frac{4\pi f_{c}R_{0}}{c}}e^{\displaystyle j\pi\frac{f^{2}_{\eta}}{K_{a}}},$ (3)
in which $K_{a}=\frac{2v^{2}}{\lambda R_{0}}$. The final step for image formation in the Range-Doppler algorithm, is to compress the data in the azimuth direction which upon performing this stage the compressed signal in the range and azimuth directions is given as [1]
$s_{\rm rcac}(t,\eta)=\sigma\;p_{r}(t-\frac{2R_{0}}{c})p_{a}(\eta)e^{\displaystyle-j\frac{4\pi f_{c}R_{0}}{c}}\times e^{\displaystyle j2\pi f_{\eta_{c}}\eta},$ (4)
where $f_{\eta_{c}}$ is the Doppler centroid frequency [1].
@@ III Statistical Modeling
In this section, we attempt to model the amplitude of the reflected signal from sea clutter. The main candidate is the Weibull distribution which its probability density functions is described as [11]
@@
The other possible options, for consideration to model the sea clutter, are the Log-normal, Inverse Gaussian, Gamma, and Rayleigh distributions with the following probability density functions [11]
@@
$p_{R}(x|\sigma)=(x/\sigma^{2})e^{\displaystyle-\frac{x^{2}}{2\sigma^{2}}},\;\;\;x\geq 0,\;\;\sigma>0.$ (9)
Based on the given probability of false alarm ($p_{\rm fa}$), we can calculate the adaptive threshold $T_{a}$ as [4]
$p_{\rm fa}=p(X>T_{a}|H_{0})=\int_{T_{a}}^{\infty}p(X|H_{0})\;dX,$ (10)
where $X$ describes the statistics of the cell under test and $T_{a}$ is the adaptive threshold. Furthermore, $H_{0}$ represents the null hypothesis. By using the experimental data that we present in this paper, we will demonstrate that among the proposed distributions, the Weibull distribution is the best distribution to model the sea clutter. Therefore, the calculation of the adaptive threshold is only performed for the Weibull distribution. Consequently, based on (10), the adaptive threshold $T_{\rm aW}$ for the Weibull distribution, which has been given in (5), is calculated as
@@
in which $x_{i}$ is the $i^{\rm th}$ realization of the random variable $x$. In this paper, we implement the 2D CFAR [4] algorithm. Subsequently, the decision making process for the cell under test is expressed as [6, 7]
@@
where $X_{\rm CUT}$ is the amplitude of the cell under test. Furthermore, $\mu_{c}$ and $\sigma_{c}$ are the sample mean and standard deviation computed from the clutter data of the local background. Moreover, $Q=\frac{T_{\rm aW}}{\hat{\mu}}$ is the detector design parameter which defines the $p_{\rm fa}$ and is set empirically in which $T_{\rm aW}$ is the adaptive threshold which has been described in (11) and $\hat{\mu}$ denotes the mean value estimated from the underlying clutter model. In addition, the $H_{0}$ and $H_{1}$ are the null and alternative hypotheses, respectively [13, 4]. The null hypothesis, $H_{0}$, represents the case in which the cell under test has been occupied by clutter only and the alternative hypothesis, $H_{1}$, on the other hand, is for the case in which the target is present and the cell under test contains the energy of the target. In the next section, we present the experimental results.
@@ IV Experimental Results
In this section, we present the result of SAR image reconstruction based on the experimental data gathered from the Canadian RADARSAT-1 satellite. The data is from Vancouver Canada. The specifications of the RADARSAT-1 satellite are given in Table.I.
@@
FM rate of radar pulse (${\rm MHz}/\mu s$) $\beta$ $0.72135$
Chirp duration ($\mu s$) $\rm T$ $41.75$
@@
We select the data related to English Bay in Vancouver, Canada. The reason for this choice is that, there are several ships in this scene that can play the role of isolated strong reflectors which will assist with observing the effect of the RCM clearly and will also be utilized later to perform the Doppler centroid frequency estimation. Furthermore, the ships in the sea water are the main subject of the paper. In other words, we attempt to detect the ships in sea clutter which is the main goal of the paper. In order to reconstruct the image, we apply the Range-Doppler algorithm to the raw data. Fig. 1-(a) shows the range compressed data based on (2). The range compressed energy of the ships can be seen as several skewed vertical lines. The skew demonstrates the effect of the RCM. In order to perform the RCM compensation we need to estimate the unambiguous Doppler centroid frequency.
@@
When the Doppler centroid frequency is larger than the PRF of the RADAR, there will be an ambiguity in the Doppler centroid frequency estimation. One way to estimate the unambiguous value for the Doppler centroid frequency is by analyzing the trajectory of the strong isolated targets. Fig. 1 shows the range compressed image of several ships in sea water. The skew in the trajectory of these targets is due to the non-zero Doppler centroid frequency. The energy of each of these targets is spread over several different range cells. The slope can easily be calculated and is equal to $0.034$ range samples per azimuth samples. To estimate the Doppler centroid frequency, the first step is to multiply the slope by $\rm\frac{c}{2Fr}$ to convert the slope from range samples to range distance and then multiply it by $\rm PRF$ to convert it from azimuth samples to azimuth time. Hence, we have $\frac{dR(\eta)}{d\eta}=198.23\;\rm m/s$. As a result, from $f_{\rm dc}=-\frac{2}{\lambda}\frac{dR(\eta)}{d\eta}$ [1], the Doppler centroid frequency can be calculated as $f_{\rm dc}=-7009\;\rm Hz$.
@@
Next, we estimate the fractional part of the Doppler centroid frequency which is $f^{\prime}_{\rm dc}$ [1]. The fractional part is essential in focusing the energy of the targets in the azimuth direction. Fig. 2 illustrates the power spectrum of the data in azimuth direction versus slow-time frequency component $f_{\eta}$. In order to reduce the effect of the noise, we have performed an averaging over the power spectrum of $2048$ range cells. In order to be able to calculate the frequency component at which the signal reaches its maximum value, we have performed curve fitting. From Fig. 2, we can estimate the fractional part of the Doppler centroid frequency as $f^{\prime}_{\rm dc}=531\;\rm Hz$. Finally, at the last stage, we perform the azimuth localization based on (4). Consequently, the reconstructed image based on the Range-Doppler algorithm is obtained which has been presented in Fig. 3.
@@
However, the image presented in Fig. 3 suffers from speckle noise [1, 14, 15]. In order to remove the effect of the speckle noise, we introduce a 2D $m\times n$ filter and slide it over the reconstructed image while solving the following optimization problem,
@@
where $a_{\rm ij}$ is the value for the $(\rm ij)^{\rm th}$ pixel and $a$ is the value selected by the optimization problem for the $(\lfloor\frac{n-1}{2}\rfloor+1,\;\lfloor\frac{m-1}{2}\rfloor+1)^{\rm th}$ pixel. In fact, by solving the optimization problem in (14) we apply a 2D median filter to the image. In other words, we replace the value of each pixel with the median of the neighboring pixels. This is proved to be a powerful method to decrease the effect of the speckle noise which, as a result, the fine structures of the image can be revealed. It should be noted that, the common practice to alleviate the effect of the speckle noise in SAR images is multi-look processing [1]. However, multi-look processing sacrifices the azimuth resolution in order to remove the effect of the speckle noise. In contrast, the median filtering approach, which we have proposed, can efficiently remove the speckle noise while it leaves the azimuth resolution intact. Furthermore, the median filtering does not smear the edges in the image which is considered to be a significant property.
@@
Fig. 4 shows the result of speckle noise reduction for the reconstructed image depicted in Fig. 3. To remove the effect of the speckle noise, we have applied the filter, given in (14), with $m=n=6$. In order to analyze the effect of the sea clutter on target detection, we focus on the specific part of the reconstructed image shown in Fig. 3 which contains several ships in sea water. The related part of the data-set has been highlighted as the area inside the green rectangle in Fig. 3. The area inside the red rectangle contains the energy from sea water and is utilized to analyze the statistical properties of the sea clutter. Fig. 5 shows the histogram for part of the image related to the area which has been located inside the red rectangle. The estimated probability density functions for the Weibull, Log-normal, Inverse Gaussian, Gamma, and Rayleigh distributions are based on (5), (6), (7), (8), and (9), respectively.
@@
Figure 5: The histogram of the experimental data inside the red box which is related to the sea clutter as well as the estimated probability density functions based on (5), (6), (7), (8), and (9). The II and $Q$ represent the imaginary and quadratic part of the estimated value for the complex radar cross section.
The estimated parameters for the given probability density functions are presented in Table.II.
@@
From Table.III, it is evident that compared to the other probability distributions, the Weibull distribution, with the lowest KL distance, can model the statistical properties of the sea clutter with higher accuracy. Therefore, we only consider the Weibull distribution for the CFAR-based analysis of target detection in sea clutter. Fig. 6 illustrates the result of applying the 2D CFAR algorithm to the part of the data-set inside the green rectangle in Fig. 3. The result presented in Fig. 6-(b) is remarkable in a sense that, by modeling the sea clutter based on the Weibull distribution and by exploiting the 2D CFAR technique we have been able to detect the ships in sea clutter with high accuracy.
Refer to captionAlong-Track [Samples]Slant-Range [Samples](a)Refer to captionAlong-Track [Samples]Slant-Range [Samples](b)
@@
To achieve the result which has been displayed in Fig. 6-(b), we have set the number of guard cells for the upper and lower wings to 60 and for the left and right wings to 90. We have also selected the number of training cells for the upper, lower, left, and right wings to be equal to 5. The probability of false alarm has been selected as $p_{\rm fa}=10^{-6}$. The Weibull parameters are estimated as $\alpha=1.9521$ and $\beta=0.4835$. Consequently, based on (11), the adaptive threshold is calculated as $T_{\rm aW}=1.856$. Moreover, the sample mean and standard deviation for the background clutter, which have been described in (13), are estimated as $\mu_{c}=0.4286$ and $\sigma_{c}=0.2294$. We consider another case in which one of the ships, as has been highlighted as the area inside the yellow rectangle in Fig. 3, is buried in the sea clutter. Moreover, it should be noted that, the image presented in Fig. 3 is contaminated with the speckle noise. In Fig. 7, we have presented the reconstructed image from a single ship which has been highlighted as the area inside the yellow box in Fig. 3.
Refer to captionAlong-Track [Samples]Slant-Range [Samples](a)Refer to captionAlong-Track [Samples]Slant-Range [Samples](b)
@@
The part of the image, shown in Fig. 7-(b), which has been highlighted as the area inside the red rectangle is utilized to estimate the histogram of the sea clutter and the result has been illustrated in Fig. 8-(a).
@@
Figure 8: a) the histogram of the part of the reconstructed image shown in Fig. 7-(b) inside the red rectangle as a representation of the statistical properties of the sea clutter. The estimated probability density function is based on the Weibull Distribution, b) the result of applying the 2D CFAR method to the image of the ship in sea clutter which has been illustrated in Fig. 7-(b)
As it is clear from Fig. 8-(a), the Weibull distribution can model the histogram with high accuracy. Next, we apply the 2D CFAR technique to the reconstructed image presented in Fig. 7-(b). We have presented this remarkable result in Fig. 8-(b). To achieve the result presented in Fig. 8-(b), we have set the number of guard cells for the upper and lower wings to 35 and for the left and right wings to 60. We have also selected the number of training cells for the upper, lower, left, and right wings to be equal to 8. The probability of false alarm has been set to $p_{\rm fa}=10^{-6}$. The Weibull parameters are estimated as $\alpha=1.9912$ and $\beta=0.2841$. Consequently, based on (11), the adaptive threshold is calculated as $T_{\rm aW}=1.0621$. Furthermore, the sample mean and standard deviation for the background clutter, which have been utilized in (13), are estimated as $\mu_{c}=0.2518$ and $\sigma_{c}=0.1321$, respectively.